OpenStudy (anonymous):

what are all the horizantal asymtotes of the graph of f(x) = (5x)/((x^2 + 1)^(1/2))

OpenStudy (anonymous):

@ganeshie8

OpenStudy (irishboy123):

\[\frac{5x}{\sqrt{x^{2\ }+\ 1}}\] \[=\pm \frac{5}{\sqrt{1+\ \frac{1}{x^2}}}\]

OpenStudy (irishboy123):